# NeuralFuse: Learning to Recover the Accuracy of Access-Limited Neural Network Inference in Low-Voltage Regimes

### v2

@@ G.3 CIFAR-100
In Tables 19 and 20, we show results on CIFAR-100 with NeuralFuse trained at 1% and 0.5% B.E.R., respectively.
@@
S.M. T.M. B.E.R. C.A. P.A. ConvL (1%) UNetL (1%) C.A. (NF) P.A. (NF) R.P. C.A. (NF) P.A. (NF) R.P. ResNet18 ResNet18 0.5% 73.7 20.9 7.4 54.8 35.8 5.2 14.9 50.6 39.3 2.8 18.4 0.35% 31.4 7.6 41.7 3.7 10.3 43.3 1.4 11.9 ResNet50 1% 73.5 3.0 1.8 44.9 2.2 2.0 -0.8 41.5 2.4 1.9 -0.6 0.5% 21.3 7.0 15.9 8.2 -5.4 17.1 7.1 -4.2 0.35% 35.7 8.6 23.7 7.1 -12.0 26.2 5.6 -9.5 VGG11 1% 64.8 8.2 5.7 41.2 9.8 5.6 1.6 37.5 10.2 5.1 2.0 0.5% 23.9 9.4 24.2 5.9 0.3 24.5 4.7 0.6 0.35% 31.3 10.0 29.0 5.4 -2.3 28.2 4.5 -3.1 VGG16 1% 67.8 7.0 3.5 44.0 7.9 3.7 0.9 39.5 10.1 4.5 3.1 0.5% 22.4 7.0 22.4 7.6 0.0 26.3 5.3 3.9 0.35% 31.1 7.2 28.1 5.9 -3.0 30.6 3.6 -0.5 VGG19 1% 67.8 10.6 4.3 44.2 13.5 6.1 2.9 40.7 15.6 6.2 5.0 0.5% 34.0 9.6 27.9 4.8 -6.1 29.3 4.6 -4.7 0.35% 42.1 9.4 33.2 48 -8.9 32.8 3.9 -9.3 VGG19 ResNet18 1% 73.7 4.6 2.9 55.5 5.8 3.7 1.2 57.3 6.8 4.4 2.2 0.5% 20.9 7.4 24.6 6.3 3.7 28.1 5.9 7.2 0.35% 31.4 7.6 31.1 5.0 -0.3 36.4 4.5 5.0 ResNet50 1% 73.5 3.0 1.8 56.1 2.8 2.1 -0.2 56.1 3.7 2.4 0.7 0.5% 21.3 7.0 18.9 8.6 -2.4 22.8 8.5 1.5 0.35% 35.7 8.6 28.7 8.2 -7.0 33.7 7.0 -2.0 VGG11 1% 64.8 8.2 5.7 52.8 12.3 8.4 4.1 53.9 15.4 9.4 7.2 0.5% 23.9 9.4 30.0 9.3 6.1 33.3 7.2 9.4 0.35% 31.3 10.0 36.5 7.7 5.2 38.8 6.5 7.5 VGG16 1% 67.8 7.0 3.5 53.6 11.2 4.4 4.2 55.2 13.6 5.2 6.6 0.5% 22.4 7.0 32.4 7.3 10.0 35.9 6.2 13.5 0.35% 31.1 7.2 39.4 6.3 8.3 42.4 4.9 11.3 VGG19 0.5% 67.8 34.0 9.6 59.4 50.2 3.1 16.2 58.7 49.1 3.5 15.1 0.35% 42.1 9.4 53.1 2.3 11.0 52.0 3.1 9.9 • [Note] S.M.: source model, used for training generators, T.M.: target model, used for testing generators, B.E.R.: the bit error rate of the target model, C.A. (%): clean accuracy, P.A. (%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A. (NF) v.s. P.A.
@@
S.M. T.M. B.E.R. C.A. P.A. ConvL (0.5%) UNetL (0.5%) C.A. (NF) P.A. (NF) R.P. C.A. (NF) P.A. (NF) R.P. ResNet18 ResNet18 0.35% 73.7 31.4 7.6 65.2 47.7 4.9 16.3 66.2 49.2 4.1 17.8 ResNet50 0.5% 73.5 21.3 7.0 62.5 24.0 9.9 2.8 63.5 26.4 9.1 5.1 0.35% 35.7 8.6 36.3 8.9 0.6 39.4 8.1 3.7 VGG11 0.5% 64.8 23.9 9.4 59.2 33.0 9.8 9.2 61.1 34.2 9.8 10.3 0.35% 31.3 10.0 40.4 8.7 9.1 41.4 9.0 10.1 VGG16 0.5% 67.8 22.4 7.0 59.5 34.7 8.0 12.3 61.4 37.5 6.8 15.2 0.35% 31.1 7.2 42.9 6.0 11.8 45.3 4.9 14.2 VGG19 0.5% 67.8 34.0 9.6 61.6 43.7 6.2 9.6 62.0 45.0 6.3 11.0 0.35% 42.1 9.4 49.0 5.5 6.8 50.5 5.3 8.3 VGG19 ResNet18 0.5% 73.7 20.9 7.4 66.1 24.9 6.7 4.0 67.8 27.7 6.8 6.8 0.35% 31.4 7.6 34.4 5.4 3.0 38.1 5.6 6.7 ResNet50 0.5% 73.5 21.3 7.0 66.2 22.7 7.8 1.4 66.7 25.4 8.0 4.2 0.35% 35.7 8.6 35.5 7.7 -0.2 38.8 7.5 3.2 VGG11 0.5% 64.8 23.9 9.4 59.9 29.3 10 5.4 61.0 31.2 9.8 7.4 0.35% 31.3 10.0 36.6 9.5 5.3 38.1 9.0 6.8 VGG16 0.5% 67.8 22.4 7.0 62.5 30.8 7.3 8.4 62.6 33.0 7.3 10.7 0.35% 31.1 7.2 40.0 6.5 8.9 42.5 5.9 11.3 VGG19 0.35% 67.8 42.1 9.4 65.6 52.0 6.2 9.8 65.5 52.6 6.1 10.4 • [Note] S.M.: source model, used for training generators, T.M.: target model, used for testing generators, B.E.R.: the bit error rate of the target model, C.A. (%): clean accuracy, P.A. (%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A. (NF) v.s. P.A.
@@ G.4 Generator Ensembling
S.M. T.M. B.E.R. C.A. P.A. ConvL (1.5%) UNetL (1.5%) C.A. (NF) P.A. (NF) R.P. C.A. (NF) P.A. (NF) R.P. ResNet18 + VGG19 ResNet18 1% 92.6 38.9 12.4 89.4 88.1 1.0 49.2 86.3 85.4 0.5 46.5 0.5% 70.1 11.6 89.2 0.2 19.1 86.1 0.2 16.0 ResNet50 1% 92.6 26.1 9.4 89.3 44.0 22 17.9 86.1 50.9 20 24.8 0.5% 61.0 10.3 80.3 6.7 19.3 78.6 3.9 17.6 VGG11 1% 88.4 42.2 11.6 89.1 77.0 5.6 34.8 85.9 82.3 4.1 40.1 0.5% 63.6 9.3 87.5 1.6 23.9 85.0 0.6 21.4 VGG16 1% 90.3 35.7 7.9 89.1 80.5 8.6 44.8 85.7 81.4 5.5 45.7 0.5% 66.6 8.1 88.2 0.7 21.6 85.0 0.7 18.4 VGG19 1% 90.5 36.0 12.0 89.2 75.1 17 39.1 86.1 83.0 3.4 47.0 0.5% 64.2 12.4 89.0 0.2 24.8 85.9 0.4 21.7 • [Note] S.M.: source model, used for training generators, T.M.: target model, used for testing generators, B.E.R.: the bit error rate of the target model, C.A. (%): clean accuracy, P.A. (%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A. (NF) v.s. P.A.
@@ Appendix H NeuralFuse on Reduced-precision Quantization and Random Bit Errors
As mentioned in Sec. 4.6, we explore the robustness of NeuralFuse to low-precision quantization of model weights and consider the case of random bit errors. Here, we demonstrate that NeuralFuse can recover not only the accuracy drop due to reduced precision, but also the drop caused by low-voltage-induced bit errors (0.5% B.E.R.) under low precision. We selected two NeuralFuse generators (ConvL and UNetL) for our experiments, and these generators were trained with the corresponding base models (ResNet18 and VGG19) at 1% B.E.R. (CIFAR-10, GTSRB) and 0.5% B.E.R. (ImageNet-10). The experimental results are shown as follows: CIFAR-10 (Sec. H.1), GTSRB (Sec. H.2), and ImageNet-10 (Sec. H.3). Similarly, for ease of comparison, we visualize the experimental results in the figures below each table. Our results show that NeuralFuse can consistently perform well in low-precision regimes as well as recover the low-voltage-induced accuracy drop.
@@ H.1 CIFAR-10
Base #Bits C.A. P.A. ConvL (1%) UNetL (1%) Model C.A. (NF) P.A. (NF) R.P. C.A. (NF) P.A. (NF) R.P. ResNet18 8 92.6 70.1 11.6 89.8 89.5 0.2 19.4 86.6 86.2 0.3 16.1 7 92.5 68.8 10.4 89.8 89.5 1.7 20.7 86.5 86.0 0.5 17.2 6 92.6 68.4 11.2 89.7 89.5 0.2 21.1 86.6 85.9 0.3 17.5 5 92.4 52.7 14.1 89.7 90.0 0.7 37.3 86.5 85.5 0.8 32.8 4 91.8 26.3 12.7 89.8 58.7 24.5 32.4 86.6 64.9 22.5 38.6 3 84.8 11.3 1.8 89.8 12.8 5.8 1.5 86.0 14.8 10.0 3.5 2 10.0 10.0 0.0 10.0 10.0 0.0 0.0 10.0 10.0 0.0 0.0 VGG19 8 90.5 64.2 12.4 89.8 89.6 8.7 25.4 87.4 86.8 0.4 22.6 7 90.3 66.5 8.5 89.8 89.6 0.2 23.1 87.4 86.7 0.3 20.2 6 90.1 59.8 13.2 89.9 89.4 3.8 29.6 87.4 86.4 0.7 26.6 5 90.2 37.7 14.1 89.8 78.0 15.8 40.3 87.2 79.8 0.8 42.1 4 87.5 14.7 6.0 89.8 27.8 18.9 13.1 87.2 34.4 20.5 19.7 3 78.3 10.5 1.5 89.7 10.9 2.6 0.4 86.8 11.0 2.9 0.5 2 10.0 10.0 0.0 10.0 10.0 0.0 0.0 10.0 10.0 0.0 0.0 • [Note] C.A. (%): clean accuracy, P.A. (%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A. (NF) v.s. P.A.
@@ H.2 GTSRB
Base #Bits C.A. P.A. ConvL (1%) UNetL (1%) Model C.A. (NF) P.A. (NF) R.P. C.A. (NF) P.A. (NF) R.P. ResNet18 8 95.5 75.2 12.7 95.7 95.3 0.5 20.1 96.2 95.7 0.3 20.5 7 95.5 69.5 10.6 95.7 95.3 0.3 25.8 96.2 95.9 0.3 26.4 6 95.4 67.2 14.4 95.7 95.2 0.5 28.0 96.2 95.7 0.5 28.5 5 95.4 48.6 18.2 95.8 92.6 5.1 44.0 96.2 94.8 2.5 46.2 4 92.6 24.6 9.8 95.9 75.6 16.2 51.0 96.2 86.6 9.5 62.0 3 67.7 5.3 3.5 95.4 18.4 15.3 13.1 96.2 25.3 22.5 20.0 2 3.8 3.8 0.0 4.1 3.8 0.0 0.0 3.8 3.8 0.0 0.0 VGG19 8 95.5 69.1 11.1 96.0 94.0 2.2 24.9 95.4 93.9 2.1 24.8 7 95.6 66.1 14.8 96.0 92.2 5.7 26.1 95.4 92.6 3.7 26.5 6 95.3 64.2 8.4 96.0 92.2 5.7 28.0 95.4 92.3 2.3 28.1 5 95.2 48.2 14.0 96.0 92.2 5.7 44.0 95.4 86.2 8.4 38.0 4 92.0 18.2 14.3 93.0 92.2 5.7 74.0 95.0 49.6 22.8 31.4 3 60.0 2.0 0.9 87.3 92.2 5.7 90.2 87.2 1.7 0.9 -0.3 2 5.9 3.8 0.0 5.9 3.8 0.0 0.0 5.9 3.8 0.0 0.0 • [Note] C.A. (%): clean accuracy, P.A. (%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A. (NF) v.s. P.A.
@@ H.3 ImageNet-10
Base #Bits C.A. P.A. ConvL (0.5%) UNetL (0.5%) Model C.A. (NF) P.A. (NF) R.P. C.A. (NF) P.A. (NF) R.P. ResNet18 8 92.2 72.3 7.0 94.0 88.0 2.0 15.7 94.0 88.1 1.4 15.8 7 92.4 70.6 13.0 94.2 86.7 4.1 16.1 93.6 87.8 3.5 17.2 6 92.4 68.9 9.9 94.2 85.1 4.8 16.2 93.6 86.4 3.7 17.5 5 91.0 60.9 13.0 94.2 82.5 6.8 21.6 94.0 83.2 5.9 22.3 4 91.4 47.4 9.8 93.8 68.6 9.8 21.2 92.6 68.7 9.2 21.3 3 85.2 28.8 11.8 89.2 44.1 14.0 15.3 89.4 42.7 14.2 13.9 2 10.0 10.0 0.0 10.0 10.0 0.0 0.0 10.0 10.0 0.0 0.0 VGG19 8 92.4 37.2 11.0 91.4 75.5 8.8 38.3 89.4 77.9 6.1 40.7 7 92.0 27.3 6.6 91.2 59.3 13.0 32.0 89.4 65.4 10.0 38.1 6 92.4 27.9 6.4 91.0 59.7 11.8 31.8 89.4 64.9 9.9 37.0 5 92.0 15.1 4.4 91.6 23.1 0.7 8.0 89.0 27.9 8.8 12.8 4 89.4 12.2 2.7 90.8 14.0 4.3 1.8 89.6 14.6 4.9 2.4 3 46.8 9.9 0.5 83.2 10.4 0.6 0.5 84.2 9.9 0.7 0.0 2 10.0 10.0 0.0 10.0 10.0 0.0 0.0 10.0 10.0 0.0 0.0 • [Note] C.A. (%): clean accuracy, P.A. (%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A. (NF) v.s. P.A.

### v3

@@ G.3 CIFAR-100
In Tables 19 and 20, we show results on CIFAR-100 with NeuralFuse trained at 1% and 0.5% BER, respectively.
@@
SM TM BER CA PA ConvL (1%) UNetL (1%) CA (NF) PA (NF) RP CA (NF) PA (NF) RP ResNet18 ResNet18 0.5% 73.7 20.9 7.4 54.8 35.8 5.2 14.9 50.6 39.3 2.8 18.4 0.35% 31.4 7.6 41.7 3.7 10.3 43.3 1.4 11.9 ResNet50 1% 73.5 3.0 1.8 44.9 2.2 2.0 -0.8 41.5 2.4 1.9 -0.6 0.5% 21.3 7.0 15.9 8.2 -5.4 17.1 7.1 -4.2 0.35% 35.7 8.6 23.7 7.1 -12.0 26.2 5.6 -9.5 VGG11 1% 64.8 8.2 5.7 41.2 9.8 5.6 1.6 37.5 10.2 5.1 2.0 0.5% 23.9 9.4 24.2 5.9 0.3 24.5 4.7 0.6 0.35% 31.3 10.0 29.0 5.4 -2.3 28.2 4.5 -3.1 VGG16 1% 67.8 7.0 3.5 44.0 7.9 3.7 0.9 39.5 10.1 4.5 3.1 0.5% 22.4 7.0 22.4 7.6 0.0 26.3 5.3 3.9 0.35% 31.1 7.2 28.1 5.9 -3.0 30.6 3.6 -0.5 VGG19 1% 67.8 10.6 4.3 44.2 13.5 6.1 2.9 40.7 15.6 6.2 5.0 0.5% 34.0 9.6 27.9 4.8 -6.1 29.3 4.6 -4.7 0.35% 42.1 9.4 33.2 48 -8.9 32.8 3.9 -9.3 VGG19 ResNet18 1% 73.7 4.6 2.9 55.5 5.8 3.7 1.2 57.3 6.8 4.4 2.2 0.5% 20.9 7.4 24.6 6.3 3.7 28.1 5.9 7.2 0.35% 31.4 7.6 31.1 5.0 -0.3 36.4 4.5 5.0 ResNet50 1% 73.5 3.0 1.8 56.1 2.8 2.1 -0.2 56.1 3.7 2.4 0.7 0.5% 21.3 7.0 18.9 8.6 -2.4 22.8 8.5 1.5 0.35% 35.7 8.6 28.7 8.2 -7.0 33.7 7.0 -2.0 VGG11 1% 64.8 8.2 5.7 52.8 12.3 8.4 4.1 53.9 15.4 9.4 7.2 0.5% 23.9 9.4 30.0 9.3 6.1 33.3 7.2 9.4 0.35% 31.3 10.0 36.5 7.7 5.2 38.8 6.5 7.5 VGG16 1% 67.8 7.0 3.5 53.6 11.2 4.4 4.2 55.2 13.6 5.2 6.6 0.5% 22.4 7.0 32.4 7.3 10.0 35.9 6.2 13.5 0.35% 31.1 7.2 39.4 6.3 8.3 42.4 4.9 11.3 VGG19 0.5% 67.8 34.0 9.6 59.4 50.2 3.1 16.2 58.7 49.1 3.5 15.1 0.35% 42.1 9.4 53.1 2.3 11.0 52.0 3.1 9.9 • Note. SM: source model, used for training generators; TM: target model, used for testing generators; BER: the bit-error rate of the target model; CA (): clean accuracy; PA (): perturbed accuracy; NF: NeuralFuse; and RP: total recovery percentage of PA (NF) vs. PA
@@
SM TM BER CA PA ConvL (0.5%) UNetL (0.5%) CA (NF) PA (NF) RP CA (NF) PA (NF) RP ResNet18 ResNet18 0.35% 73.7 31.4 7.6 65.2 47.7 4.9 16.3 66.2 49.2 4.1 17.8 ResNet50 0.5% 73.5 21.3 7.0 62.5 24.0 9.9 2.8 63.5 26.4 9.1 5.1 0.35% 35.7 8.6 36.3 8.9 0.6 39.4 8.1 3.7 VGG11 0.5% 64.8 23.9 9.4 59.2 33.0 9.8 9.2 61.1 34.2 9.8 10.3 0.35% 31.3 10.0 40.4 8.7 9.1 41.4 9.0 10.1 VGG16 0.5% 67.8 22.4 7.0 59.5 34.7 8.0 12.3 61.4 37.5 6.8 15.2 0.35% 31.1 7.2 42.9 6.0 11.8 45.3 4.9 14.2 VGG19 0.5% 67.8 34.0 9.6 61.6 43.7 6.2 9.6 62.0 45.0 6.3 11.0 0.35% 42.1 9.4 49.0 5.5 6.8 50.5 5.3 8.3 VGG19 ResNet18 0.5% 73.7 20.9 7.4 66.1 24.9 6.7 4.0 67.8 27.7 6.8 6.8 0.35% 31.4 7.6 34.4 5.4 3.0 38.1 5.6 6.7 ResNet50 0.5% 73.5 21.3 7.0 66.2 22.7 7.8 1.4 66.7 25.4 8.0 4.2 0.35% 35.7 8.6 35.5 7.7 -0.2 38.8 7.5 3.2 VGG11 0.5% 64.8 23.9 9.4 59.9 29.3 10 5.4 61.0 31.2 9.8 7.4 0.35% 31.3 10.0 36.6 9.5 5.3 38.1 9.0 6.8 VGG16 0.5% 67.8 22.4 7.0 62.5 30.8 7.3 8.4 62.6 33.0 7.3 10.7 0.35% 31.1 7.2 40.0 6.5 8.9 42.5 5.9 11.3 VGG19 0.35% 67.8 42.1 9.4 65.6 52.0 6.2 9.8 65.5 52.6 6.1 10.4 • Note. SM: source model, used for training generators; TM: target model, used for testing generators; BER: the bit-error rate of the target model; CA (): clean accuracy; PA (): perturbed accuracy; NF: NeuralFuse; and RP: total recovery percentage of PA (NF) vs. PA
@@ G.4 Generator Ensembling
SM TM BER CA PA ConvL (1.5%) UNetL (1.5%) CA (NF) PA (NF) RP CA (NF) PA (NF) RP ResNet18 + VGG19 ResNet18 1% 92.6 38.9 12.4 89.4 88.1 1.0 49.2 86.3 85.4 0.5 46.5 0.5% 70.1 11.6 89.2 0.2 19.1 86.1 0.2 16.0 ResNet50 1% 92.6 26.1 9.4 89.3 44.0 22 17.9 86.1 50.9 20 24.8 0.5% 61.0 10.3 80.3 6.7 19.3 78.6 3.9 17.6 VGG11 1% 88.4 42.2 11.6 89.1 77.0 5.6 34.8 85.9 82.3 4.1 40.1 0.5% 63.6 9.3 87.5 1.6 23.9 85.0 0.6 21.4 VGG16 1% 90.3 35.7 7.9 89.1 80.5 8.6 44.8 85.7 81.4 5.5 45.7 0.5% 66.6 8.1 88.2 0.7 21.6 85.0 0.7 18.4 VGG19 1% 90.5 36.0 12.0 89.2 75.1 17 39.1 86.1 83.0 3.4 47.0 0.5% 64.2 12.4 89.0 0.2 24.8 85.9 0.4 21.7 • Note. SM: source model, used for training generators; TM: target model, used for testing generators; BER: the bit-error rate of the target model; CA (): clean accuracy; PA (): perturbed accuracy; NF: NeuralFuse; and RP: total recovery percentage of PA (NF) vs. PA
@@ Appendix H NeuralFuse on Reduced-precision Quantization and Random Bit Errors
As mentioned in Sec. 4.6, we explore the robustness of NeuralFuse to low-precision quantization of model weights and consider the case of random bit errors. Here, we demonstrate that NeuralFuse can recover not only the accuracy drop due to reduced precision, but also the drop caused by low-voltage-induced bit errors (0.5% BER) under low precision. We selected two NeuralFuse generators (ConvL and UNetL) for our experiments, and these generators were trained with the corresponding base models (ResNet18 and VGG19) at 1% BER (CIFAR-10, GTSRB) and 0.5% BER (ImageNet-10). The experimental results are shown as follows: CIFAR-10 (Sec. H.1), GTSRB (Sec. H.2), and ImageNet-10 (Sec. H.3). Similarly, for ease of comparison, we visualize the experimental results in the figures below each table. Our results show that NeuralFuse can consistently perform well in low-precision regimes as well as recover the low-voltage-induced accuracy drop.
@@ H.1 CIFAR-10
Base #Bits CA PA ConvL (1%) UNetL (1%) Model CA (NF) PA (NF) RP CA (NF) PA (NF) RP ResNet18 8 92.6 70.1 11.6 89.8 89.5 0.2 19.4 86.6 86.2 0.3 16.1 7 92.5 68.8 10.4 89.8 89.5 1.7 20.7 86.5 86.0 0.5 17.2 6 92.6 68.4 11.2 89.7 89.5 0.2 21.1 86.6 85.9 0.3 17.5 5 92.4 52.7 14.1 89.7 90.0 0.7 37.3 86.5 85.5 0.8 32.8 4 91.8 26.3 12.7 89.8 58.7 24.5 32.4 86.6 64.9 22.5 38.6 3 84.8 11.3 1.8 89.8 12.8 5.8 1.5 86.0 14.8 10.0 3.5 2 10.0 10.0 0.0 10.0 10.0 0.0 0.0 10.0 10.0 0.0 0.0 VGG19 8 90.5 64.2 12.4 89.8 89.6 8.7 25.4 87.4 86.8 0.4 22.6 7 90.3 66.5 8.5 89.8 89.6 0.2 23.1 87.4 86.7 0.3 20.2 6 90.1 59.8 13.2 89.9 89.4 3.8 29.6 87.4 86.4 0.7 26.6 5 90.2 37.7 14.1 89.8 78.0 15.8 40.3 87.2 79.8 0.8 42.1 4 87.5 14.7 6.0 89.8 27.8 18.9 13.1 87.2 34.4 20.5 19.7 3 78.3 10.5 1.5 89.7 10.9 2.6 0.4 86.8 11.0 2.9 0.5 2 10.0 10.0 0.0 10.0 10.0 0.0 0.0 10.0 10.0 0.0 0.0 • Note. CA (): clean accuracy; PA (): perturbed accuracy; NF: NeuralFuse; and RP: total recovery percentage of PA (NF) vs. PA
@@ H.2 GTSRB
Base #Bits CA PA ConvL (1%) UNetL (1%) Model CA (NF) PA (NF) RP CA (NF) PA (NF) RP ResNet18 8 95.5 75.2 12.7 95.7 95.3 0.5 20.1 96.2 95.7 0.3 20.5 7 95.5 69.5 10.6 95.7 95.3 0.3 25.8 96.2 95.9 0.3 26.4 6 95.4 67.2 14.4 95.7 95.2 0.5 28.0 96.2 95.7 0.5 28.5 5 95.4 48.6 18.2 95.8 92.6 5.1 44.0 96.2 94.8 2.5 46.2 4 92.6 24.6 9.8 95.9 75.6 16.2 51.0 96.2 86.6 9.5 62.0 3 67.7 5.3 3.5 95.4 18.4 15.3 13.1 96.2 25.3 22.5 20.0 2 3.8 3.8 0.0 4.1 3.8 0.0 0.0 3.8 3.8 0.0 0.0 VGG19 8 95.5 69.1 11.1 96.0 94.0 2.2 24.9 95.4 93.9 2.1 24.8 7 95.6 66.1 14.8 96.0 92.2 5.7 26.1 95.4 92.6 3.7 26.5 6 95.3 64.2 8.4 96.0 92.2 5.7 28.0 95.4 92.3 2.3 28.1 5 95.2 48.2 14.0 96.0 92.2 5.7 44.0 95.4 86.2 8.4 38.0 4 92.0 18.2 14.3 93.0 92.2 5.7 74.0 95.0 49.6 22.8 31.4 3 60.0 2.0 0.9 87.3 92.2 5.7 90.2 87.2 1.7 0.9 -0.3 2 5.9 3.8 0.0 5.9 3.8 0.0 0.0 5.9 3.8 0.0 0.0 • Note. CA (): clean accuracy; PA (): perturbed accuracy; NF: NeuralFuse; and RP: total recovery percentage of PA (NF) vs. PA
@@ H.3 ImageNet-10
Base #Bits CA PA ConvL (0.5%) UNetL (0.5%) Model CA (NF) PA (NF) RP CA (NF) PA (NF) RP ResNet18 8 92.2 72.3 7.0 94.0 88.0 2.0 15.7 94.0 88.1 1.4 15.8 7 92.4 70.6 13.0 94.2 86.7 4.1 16.1 93.6 87.8 3.5 17.2 6 92.4 68.9 9.9 94.2 85.1 4.8 16.2 93.6 86.4 3.7 17.5 5 91.0 60.9 13.0 94.2 82.5 6.8 21.6 94.0 83.2 5.9 22.3 4 91.4 47.4 9.8 93.8 68.6 9.8 21.2 92.6 68.7 9.2 21.3 3 85.2 28.8 11.8 89.2 44.1 14.0 15.3 89.4 42.7 14.2 13.9 2 10.0 10.0 0.0 10.0 10.0 0.0 0.0 10.0 10.0 0.0 0.0 VGG19 8 92.4 37.2 11.0 91.4 75.5 8.8 38.3 89.4 77.9 6.1 40.7 7 92.0 27.3 6.6 91.2 59.3 13.0 32.0 89.4 65.4 10.0 38.1 6 92.4 27.9 6.4 91.0 59.7 11.8 31.8 89.4 64.9 9.9 37.0 5 92.0 15.1 4.4 91.6 23.1 0.7 8.0 89.0 27.9 8.8 12.8 4 89.4 12.2 2.7 90.8 14.0 4.3 1.8 89.6 14.6 4.9 2.4 3 46.8 9.9 0.5 83.2 10.4 0.6 0.5 84.2 9.9 0.7 0.0 2 10.0 10.0 0.0 10.0 10.0 0.0 0.0 10.0 10.0 0.0 0.0 • Note. CA (): clean accuracy; PA (): perturbed accuracy; NF: NeuralFuse; and RP: total recovery percentage of PA (NF) vs. PA